### v1

@@ Abstract
Transfer and Koopman operator methods offer a framework for representing complex, nonlinear dynamical systems via linear transformations, enabling for a deeper understanding of the underlying dynamics. The spectrum of these operators provide important insights into system predictability and emergent behaviour, although efficiently estimating them from data can be challenging. We tackle this issue through the lens of general operator and representational learning, in which we approximate these linear operators using efficient finite-dimensional representations. Specifically, we machine-learn orthonormal, locally supported basis functions that are dynamically tailored to the system. This learned basis provides a particularly accurate approximation of the operator’s action as well as a nearly invariant finite-dimensional subspace. We illustrate our approach with examples that showcase the retrieval of spectral properties from the estimated operator, and emphasise the dynamically adaptive quality of the machine-learned basis.
@@ 1 Introduction
Nonlinear dynamical systems emerge in a variety of domains, including biology, ecology, physics, and engineering; understanding their behaviour is critical for prediction and control. However, studying nonlinear systems is difficult due to their complex and frequently high-dimensional and chaotic behaviour. A promising approach is to study a nonlinear system via an associated linear operator acting on observables (functions of the state). In this operator-theoretic framework, the nonlinear evolution in state space is lifted to a linear (but infinite-dimensional) evolution in a function space. For example, given a discrete dynamical system $x_{n+1}=T(x_{n})$ with state space $X$, one can define the transfer operator $\mathcal{L}$ acting on a function $g:X\to\mathbb{R}$ by composition:
$\mathcal{L}g\coloneq g\circ T^{-1}$
@@
A major challenge in typical applications is that we do not know a priori which observables or subspaces will yield a good finite-dimensional representation of $\mathcal{L}$. Many traditional approaches assume a predefined set of basis functions – e.g. monomials, Fourier modes, or the library of observables supplied to Extended Dynamic Mode Decomposition (EDMD) [27] – that may not be well suited to the system at hand. For example, data-driven approximations of the Koopman operator, such as DMD [25, 2] and its variations, depend on the user supplying observables that (approximately) span an invariant subspace under $\mathcal{L}$. If key observables are omitted, important dynamics can be missed, and a lack of invariance can contribute to estimation errors. Consequently, identifying coordinate transformations or feature spaces in which a nonlinear system behaves linearly remains an active and widely researched topic [16, 1, 11, 18].
Recent research has concentrated on learning suitable observables or invariant subspaces from data rather than defining them beforehand. Takeishi et al. [26] introduced a fully data-driven method to identify a Koopman-invariant subspace by optimising over a class of candidate functions (represented by neural networks). Similarly, one can use deep learning to discover Koopman eigenfunctions by training an autoencoder whose latent variables evolve linearly in time [21]. Such an approach would yield intrinsic coordinates that linearize the dynamics.
@@
In parallel, there has been significant progress in learning operators (maps between function spaces) directly using neural networks. This is built upon the universal approximation theorem for operators [6, 17], which extends the classic universal approximation property of neural nets from finite-dimensional functions to infinite-dimensional functionals. Lu et al. proposed the Deep Operator Network (DeepONet) [20] architecture. DeepONet consists of two subnetworks (a branch network that encodes the input function via evaluations at a fixed set of sensor points, and a trunk network that encodes the location $x$ where the output is evaluated) and can learn nonlinear operator mappings. Another example is the Basis Operator Network (BasisONet) of Hua and Lu [14], which uses a neural-network-based dual autoencoder setting to learn a lower-dimensional representation (basis) for functions in the input and output function spaces. Over the past few years, different architectures have been proposed for various scenarios, including Fourier Neural Operators [19] and PCA-Net [3]. A comprehensive review of operator learning techniques may be found in [4], including insightful parallels to the matrix recovery problem.
Despite this progress, there remains a gap in jointly learning the operator and an invariant subspace tailored to that operator. In this work, we propose a neural-network architecture that simultaneously learns an approximation of the transfer operator $\mathcal{L}$ and a set of basis functions with desirable numerical properties (such as orthonormality and local support) that span a nearly invariant subspace of $L^{2}(X)$. Our learned basis automatically adapts to the dynamics to create highly efficient (in terms of basis functions required) and accurate bases for finite-dimensional representations of the transfer operator. We mainly discuss the transfer operator, but all statements also hold for the Koopman operator.
@@
An outline of the paper is as follows. Section 2 introduces the Single Autoencoder Basis Operator Network (SABON) and proves a universal-approximation theorem for the presented setting. Section 3 validates SABON, first on a circle-rotation map—showing that spectral properties of the transfer operator can be recovered from the learned basis and approximator—and second on a perturbed Arnold cat map, where the basis adapts to the anisotropic dynamics and outperforms an equal-cardinality Fourier basis. The final section discusses the main findings and sketches directions for future work.
@@ 2 Methodology
We introduce notation that will be useful throughout the paper. Let the state space $(X,\mathcal{B},\mu)$ be a probability space with $X$ compact, and let $T\colon X\to X$ a measure‑preserving transformation. We consider $\{x_{i}\}_{i=1,\dots,n}\subset X$ to be a discretization of the state space, for example a set of grid or sampling points. An observable $g\in L^{2}(X)$ is represented by its values on this grid, that is $[g(x_{1}),g(x_{2}),\ldots,g(x_{n})]^{\top}$. The action of the transfer operator produces a new map $\mathcal{L}(g)$, whose values on the grid are $\big{[}(\mathcal{L}g)(x_{1}),\ldots,(\mathcal{L}g)(x_{n})\big{]}^{\top}$. If $V\subset L^{2}(X)$ is a finite-dimensional invariant subspace spanned by basis functions ${\phi_{1},\dots,\phi_{N}}$, then the restriction of $\mathcal{L}$ to $V$ can be represented by an $N\times N$ matrix $L$ such that $\mathcal{L}\phi_{j}=\sum_{k=1}^{N}L_{kj}\phi_{k}$. If $V$ is not $\mathcal{L}$-invariant, a finite-dimensional approximation $\hat{\mathcal{L}}$ is often obtained by first applying $\mathcal{L}$ to the basis elements of $V$ and then projecting the resulting functions back into $V$. The quality of this approximation depends on the choice of subspace $V$. Two main components are the richness of the space $V$ and the error incurred due to projection into $V$. The latter error is greatly reduced if $V$ is approximately invariant under $\mathcal{L}$. Invariant (or approximately invariant) subspaces thus play a key role in accurately approximating $\mathcal{L}$.
We will adopt the standard convention of scalar feed-forward neural networks as a composition of affine maps and nonlinear activations, $\phi\colon\mathbb{R}^{d}\to\mathbb{R}$. Stacking $Q$ such networks gives a vector‑valued network $\psi\colon\mathbb{R}^{d}\to\mathbb{R}^{Q}$, written as
$\hat{\phi}(x)=\bigl{(}\phi_{1}(x),\dots,\phi_{Q}(x)\bigr{)}^{\top}$
@@ 2.1 Single Autoencoder Basis Operator Networks (SABON)
To address the problem of simultaneously learning an approximator to the Transfer Operator and a suitable invariant subspace, we propose the following architecture composed of a single autoencoder and a standard neural network.
Figure 1: SABON consists of an encoder, a latent projection, two identical decoders, and a linear map $\mathcal{G}$. The encoder learns basis functions $\{\phi_{j}\}_{j=1}^{N}$, while $\mathcal{G}\colon\mathbb{R}^{N}\to\mathbb{R}^{N}$ approximates the action of ^$\hat{\mathcal{L}}(g)$ on $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. Using two decoders - one for $g$ and one for ^$\hat{\mathcal{L}}(g)$ enforces a basis that is suitable for representing both observables and their transformed counterparts.
Two inputs are needed for the above architecture: the observable function $g$, represented by its evaluation over the sampling points $[g(x_{1}),g(x_{2}),\ldots,g(x_{n})]^{\top}$ and the discretization $\{x_{i}\}_{i=1,\dots,n}\subset X$, where all maps are evaluated. We explain the SABON components in turn.
@@
The encoder in SABON is a vector‑valued neural network
$\hat{\phi}\colon\mathbb{R}^{n}\longrightarrow\mathbb{R}^{N}$
that sends the grid vector $[g(x_{1}),\dots,g(x_{n})]^{\top}$ to the evaluations
$\bigl{[}\phi_{1}(x_{1}),\dots,\phi_{N}(x_{n})\bigr{]}^{\top}$
Its $N$ scalar components $\{\phi_{j}\}_{j=1}^{N}$ form the learned neural basis of the subspace $V$ introduced above.
@@
The projection block computes $\{\langle g,\phi_{j}\rangle\}_{j=1,\dots,N}$ for an observable $g\in L^{2}(X)$ and a set of learned basis functions $\{\phi_{j}\}_{j=1}^{N}$. In more general terms, we define $\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}:L^{2}(X)\to\mathbb{R}^{N}$ via
@@
We shall omit the subscript of $\mathcal{P}$ throughout the rest of the paper.
@@
The action of the transfer/Koopman operator is learned in the form of a linear map on block $\mathcal{G}$. The coefficients $\langle g,\phi_{j}\rangle$, $j=1,\ldots,N$ in the latent space are input to the subnetwork $\mathcal{G}$, which learns the action of $\hat{\mathcal{L}}$ in the latent space. That is, the block $\mathcal{G}$ consists of a learnable linear map capable of transforming the expanding coefficients of $g$ on $\{\phi_{j}\}_{j=1}^{N}$ into the expanding coefficients of ^$\hat{\mathcal{L}}(g)$ onto the same basis.
Decoder $\mathcal{D}$
The architecture employs two identical decoder instances: (i) the upper decoder that reconstructs the transformed observable ^$\hat{\mathcal{L}}(g)$, and (ii) the lower decoder that reconstructs the original observable $g$ shown at the bottom of the diagram. Given the latent coefficients $c_{j}:=\langle g,\phi_{j}\rangle$, the lower decoder produces
$\hat{g}(x)=\sum_{j=1}^{N}c_{j}\,\phi_{j}(x),$
which, for an orthonormal basis $\{\phi_{j}\}_{j=1}^{N}$, is the orthogonal projection of $g$ onto $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. In that same spirit, the upper decoder receives the coefficients $(\alpha_{1},\dots,\alpha_{N})$ supplied by the block $\mathcal{G}$ (defined next) and reconstructs
@@
[…]
$\mathcal{D}_{\{\phi_{j}\}_{j=1}^{N}}(a_{1},\dots,a_{N})=\sum_{j=1}^{N}a_{j}\,\phi_{j}$
To simplify the notation, we will also omit the subscript in the sequel. The coefficients serving as input to $\mathcal{D}$ will always be clear from the context (i.e. which maps are composed with $\mathcal{D}$).
Both decoders will appear in the loss term: the lower one is used in the reconstruction loss for the observable $g$, while the decoder in the top row is considered in the reconstruction loss for the transformed observable ^$\hat{\mathcal{L}}(g)$. This enforces that the learned basis $\{\phi_{j}\}_{j=1}^{N}$ is simultaneously well suited for expanding the observable itself and the action of the operator on that observable.
One of the key features of basis operator networks is that they simultaneously optimise the learned subspace and the operator action. This is performed through a suitable loss function, adapted from [14] to include a sparsity penalty that enforces local support of the basis functions. The loss function has three components (in practice these are evaluated with discretized versions):
@@
    The deviation of the autoencoder blocks from the identity are optimised with the relative $L_{2}$ loss with respect to the identity operator on $L^{2}(X)$ ($Id$).
    (2) $\mathcal{E}_{1}(\mathcal{P},\mathcal{D},g):=\frac{\|Id(g)-\mathcal{D}\circ\mathcal{P}(g)\|_{2}}{\|Id(g)\|_{2}}.$
@@
    The network $\mathcal{G}$, responsible for learning the projected action of $\mathcal{L}$ on $V$ has the loss term:
    (3) $\mathcal{E}_{2}(\mathcal{L},\mathcal{G},\mathcal{P},\mathcal{D},g):=\frac{\|\mathcal{L}(g)-\mathcal{D}\circ\mathcal{G}\circ\mathcal{P}(g)\|_{2}}{\|\mathcal{L}(g)\|_{2}}.$
  3. 3.
    The third term of the loss function is a sparsity penalty to encourage the learned basis functions to have local support
    (4) $\mathcal{E}_{3}(\{\phi_{j}\}_{j=1}^{N})=\frac{1}{N}\sum_{j=1}^{N}\|\phi_{j}\|_{L^{1}}.$
@@
(5) $J(\mathcal{L},\mathcal{G},I_{d},\mathcal{P},\mathcal{D},g)=\beta_{1}\mathcal{E}_{1}(Id,\mathcal{P},\mathcal{D},g)+\beta_{2}\mathcal{E}_{2}(\mathcal{L},\mathcal{G},\mathcal{P},\mathcal{D},g)+\beta_{3}\mathcal{E}_{3}(\{\phi_{j}\}_{j=1}^{N}),$
where $\beta_{1},\beta_{2},\beta_{3}>0$. Finally, we discuss the activation functions employed in the proposed architecture. For the $\mathcal{G}$ network, which is learning the action of a linear operator (transfer or Koopman), we selected a linear activation function, namely the identity function, and set the bias term to zero. The activation functions for the encoder block ($\mathcal{E}$) can be chosen among the usual candidates, and we had the best results with ReLU in our experiments.
The combined architecture and loss in SABON are structured so that the learned basis $\{\phi_{j}\}_{j=1}^{N}$
  • is locally supported, enforced by the sparsity term in the loss function;
  • is approximately orthogonal;
  • spans a subspace $V$ that is (approximately) invariant under the Perron–Frobenius/Koopman operator;
  • yields small projection error for any observable $g$, thanks to the loss term $\mathcal{E}_{1}$; and
  • yields equally small projection error for the transformed observable ^$\hat{\mathcal{L}}(g)$, given the term $\mathcal{E}_{2}$ in the loss function.
Thus $\mathcal{E}_{2}$ simultaneously optimises for a good finite–dimensional approximation of the operator action and enforces that $\{\phi_{j}\}_{j=1}^{N}$ spans a space $V$ well suited for representing ^$\hat{\mathcal{L}}(g)$. A final rescaling step yields an orthonormal basis.

### v2

@@ Abstract
Transfer and Koopman operator methods offer a framework for representing complex, nonlinear dynamical systems via linear transformations, enabling a deeper understanding of the underlying dynamics. The spectra of these operators provide important insights into system predictability and emergent behaviour, although efficiently estimating them from data can be challenging. We approach this issue through the lens of general operator and representational learning, in which we approximate these linear operators using efficient finite-dimensional representations. Specifically, we machine-learn orthonormal basis functions that are dynamically tailored to the system. This learned basis provides a particularly accurate approximation of the operator’s action as well as a nearly invariant finite-dimensional subspace. We illustrate our approach with examples that showcase the retrieval of spectral properties from the estimated operator, and emphasise the dynamically adaptive quality of the machine-learned basis.
@@ 1 Introduction
Nonlinear dynamical systems emerge in a variety of domains, including biology, ecology, physics, and engineering; understanding their behaviour is critical for prediction and control. However, studying nonlinear systems is difficult due to their complex and frequently high-dimensional and chaotic behaviour. A promising approach is to study a nonlinear system via an associated linear operator acting on observables (nonlinear functions of the state). In this operator-theoretic framework, the nonlinear evolution in state space is lifted to a linear (but infinite-dimensional) evolution in a function space. For example, given a discrete dynamical system $x_{n+1}=T(x_{n})$ with state space $X$, one can define the transfer operator $\mathcal{L}$ acting on a function $g:X\to\mathbb{R}$ by composition:
$\mathcal{L}g\coloneq g\circ T^{-1}.$
@@
A major challenge in typical applications is that we do not know a priori which observables or subspaces will yield a good finite-dimensional representation of $\mathcal{L}$. Many traditional approaches assume a predefined set of basis functions — e.g. monomials, Fourier modes, or the library of observables supplied to Extended Dynamic Mode Decomposition (EDMD) [34] — that may not be well suited to the system at hand. For example, data-driven approximations of the Koopman operator, such as DMD [31, 2] and its variations, depend on the user supplying observables that span a subspace that is approximately invariant under $\mathcal{L}$. If key observables are omitted, important dynamics can be missed, while a lack of invariance can contribute to estimation errors. Consequently, identifying coordinate transformations or feature spaces in which a nonlinear system behaves linearly remains an active and widely researched topic [21, 1, 13, 23].
Several approaches attempt to learn such coordinate transformations. Brunton et al. consider a sparse dictionary of functions with fixed form to obtain a finite Koopman-invariant subspace for control [7]. Starting from a specially structured dictionary, Johnson et al. [17] are able to construct invariant subspaces and, in a companion work, propose a test for their invariance [18]. Turning to the use of neural networks, Li et al. [24] employ a neural dictionary (neural networks as functions), updating the basis and Koopman matrix separately. As a result, operator errors never backpropagate into the dictionary, and the learned observables are not adapted to the underlying dynamics.
Unifying both operator and embedding learning, Otto and Rowley’s linearly recurrent autoencoder network (LRAN) forces an encoded version of the state space to evolve linearly in a Koopman-invariant subspace [29]. However, the loss is skewed towards a good reconstruction of the state space from its encoded version. The network is tuned to embed the state space compactly rather than to learn a function space that is nearly closed under $\mathcal{L}$. Other encoder-decoder schemes, such as Yeung et al. [35] and control-oriented work by [32] follow a similar approach. Takeishi et al. [33] also learn observables by minimizing a one-step residual; at the global optimum these observables span an invariant subspace. Using a modified autoencoder structure, Lusch, Brunton, and Kutz consider the spectrum of the operator and learn Koopman eigenfunctions [27], but tuning the network parameters is not straightforward.
@@
In parallel, there has been significant progress in learning nonlinear operators (nonlinear maps between function spaces) with neural networks. This work is built upon the universal approximation theorem for operators [8, 22], which extends the classic universal approximation property of neural nets from finite-dimensional functions to infinite-dimensional functionals. Lu et al. [26] have proposed the Deep Operator Network (DeepONet) architecture. DeepONet consists of two subnetworks (a branch network that encodes the input function via evaluations at a fixed set of sensor points, and a trunk network that encodes the location where the output is evaluated) and can learn nonlinear operator mappings. Another example is the Basis Operator Network (BasisONet) of Hua and Lu [16], which uses a neural-network-based dual autoencoder setting to learn a lower-dimensional representation (basis) for functions in the input and output function spaces. Over the past few years, different architectures have been proposed for various scenarios, including Fourier Neural Operators [25], and PCA-Net [3]. A comprehensive review of operator learning techniques may be found in [4], including insightful parallels to the matrix recovery problem.
In this current work we take the BasisONet architecture as a starting point, and exploit properties of Perron–Frobenius and Koopman operators to reduce the number of encoding blocks and produce a single basis of functions that span an invariant subspace. Our Single Autoencoder Basis Operator Network (SABON) learns a basis with no predefined functional form and trains it with a relative $L^{2}$ loss. For every training observable we minimise a residual measuring how well a finite matrix represents $\mathcal{L}$ on the subspace spanned by the learned basis. Because of the inner-product structure of our projection block, gradient descent methods naturally push the basis toward orthogonality, improving numerical conditioning. An optional sparsity penalty encourages local support (Example 1). The combined learning of both the operator estimate and the collection of functions produces a basis that is adapted to the underlying dynamics, which further enhances the invariance property (Example 2).
@@
An outline of the paper is as follows. Section 2 introduces the Single Autoencoder Basis Operator Network (SABON) and proves a universal-approximation theorem for the presented setting. Section 3 first validates SABON on a circle rotation map, and shows that spectral properties of the transfer operator can be recovered from the learned basis and approximator. Next, we apply SABON to a nonlinear hyperbolic map of the 2-torus, and demonstrate that the learned basis adapts to the strongly anisotropic dynamics and outperforms an equal-cardinality Fourier basis in terms of basis invariance and eigenfunction approximation. The final section discusses our main findings and sketches directions for future work.
@@ 2 Methodology
Let $X\subset\mathbb{R}^{d}$ denote a compact, metrizable space of intrinsic dimension $d^{\prime}\leq d$, where elements of $X$ are parameterised by $d$-tuples of real numbers. The set $X$ is the domain of our nonlinear dynamics, which are generated by the map $T\colon X\to X$. We assume that $\mathcal{L}$ is continuous; sufficient conditions on $T$ ensuring this continuity are given in Remark 2.2. Let $\{x_{i}\}_{i=1,\dots,n}\subset X$ be a set of sample points representing a discretization of the state space. We assume that in the large-data limit these sample points are distributed according to a probability measure $\nu$ on $X$, and in what follows we consider the space of square-integrable real-valued functions $L^{2}(X,\nu)$, which we write as $L^{2}(X)$ for brevity. An observable $g\in L^{2}(X)$ is represented by its values on this grid, that is $[g(x_{1}),g(x_{2}),\ldots,g(x_{n})]^{\top}$. The action of the transfer operator produces a new function $\mathcal{L}(g)$, whose values on the grid are $\big{[}(\mathcal{L}g)(x_{1}),\ldots,(\mathcal{L}g)(x_{n})\big{]}^{\top}$. If $V\subset L^{2}(X)$ is a finite-dimensional invariant subspace spanned by basis functions ${\phi_{1},\dots,\phi_{N}}$, then the restriction of $\mathcal{L}$ to $V$ can be represented by an $N\times N$ matrix $L$ such that $\mathcal{L}\phi_{j}=\sum_{k=1}^{N}L_{kj}\phi_{k}$. If $V$ is not $\mathcal{L}$-invariant, a finite-rank approximation $\hat{\mathcal{L}}$ is often obtained by first applying $\mathcal{L}$ to the basis elements of $V$ and then projecting the resulting functions back into $V$. The quality of this approximation depends on the choice of subspace $V$. Two important contributors to approximation accuracy are the richness of the space $V$ and the error incurred due to projection into $V$. The latter error is greatly reduced if $V$ is approximately invariant under $\mathcal{L}$. Invariant (or approximately invariant) subspaces thus play a key role in accurately approximating $\mathcal{L}$.
In preparation for the neural learning, we suppose that the state space $X$ may be embedded in $\mathbb{R}^{d}$. We will adopt the standard convention of scalar feed-forward neural networks as a composition of affine maps and nonlinear activations, $\phi\colon\mathbb{R}^{d}\to\mathbb{R}$. Stacking $N$ such networks gives a vector‑valued network $\hat{\phi}(x)\colon\mathbb{R}^{d}\to\mathbb{R}^{N}$, written as
(1) $\hat{\phi}(x)=\bigl{(}\phi_{1}(x),\dots,\phi_{N}(x)\bigr{)}^{\top}$
@@ 2.1 Single Autoencoder Basis Operator Networks (SABON)
To address the problem of simultaneously learning an approximator to the Transfer Operator and a suitable invariant subspace, we employ an architecture comprising four stages: (i) an encoder $\mathcal{E}$ that constructs basis functions evaluated on a training grid, (ii) a block $\mathcal{P}$ that projects observables onto the learned basis, (iii) a linear latent map $\mathcal{G}$ that approximates the transfer operator in the coefficient space, and (iv) a reconstructor $\mathcal{R}$ that lifts these coefficients back to the function space by expanding them in the learned basis.
Figure 1: SABON consists of an encoder $\mathcal{E}$, a latent projection $\mathcal{P}$, a linear map $\mathcal{G}$, and a reconstruction map $\mathcal{R}$. The encoder learns basis functions $\{\phi_{j}\}_{j=1}^{N}$, while $\mathcal{G}\colon\mathbb{R}^{N}\to\mathbb{R}^{N}$ approximates the action of $\mathcal{L}$ on $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$. The reconstruction map recovers a function from its coefficients in the latent space.
Two inputs are needed for the above architecture: the discretization $\{x_{i}\}_{i=1,\dots,n}\subset X$ where all maps are evaluated, and the observable function $g$, represented by its evaluation over these sampling points $[g(x_{1}),g(x_{2}),\ldots,g(x_{n})]^{\top}$.
We explain the SABON components in turn.
@@
The encoder in SABON is a stack of vector‑valued neural networks, each applied pointwise to one grid element. Recall from (1) that $\hat{\phi}\colon\mathbb{R}^{d}\longrightarrow\mathbb{R}^{N}$ is the vector-valued network defined in the embedding of $X$ in $\mathbb{R}^{d}$. Then the encoder
$\mathcal{E}\colon(\mathbb{R}^{d})^{n}\longrightarrow(\mathbb{R}^{N})^{n}\cong\mathbb{R}^{Nn},\qquad\mathcal{E}\bigl{(}(x_{1},\dots,x_{n})\bigr{)}=\bigl{(}\hat{\phi}(x_{1}),\dots,\hat{\phi}(x_{n})\bigr{)}$
acts by evaluating $\hat{\phi}$ at every grid point. Written componentwise, the output is the column-major vector
$\bigl{[}\phi_{1}(x_{1}),\dots,\phi_{N}(x_{1}),\phi_{1}(x_{2}),\dots,\phi_{N}(x_{2}),\dots,\phi_{1}(x_{n}),\dots,\phi_{N}(x_{n})\bigr{]}^{\top}\in\mathbb{R}^{Nn}.$
The collection $\{\phi_{j}\}_{j=1}^{N}$ forms the learned neural basis of the subspace $V$ introduced above.
@@
The projection block computes111in practice, this inner product is evaluated as $\frac{1}{n}\sum_{i=1}^{n}g(x_{i})\phi_{j}(x_{i})$. $\{\langle g,\phi_{j}\rangle\}_{j=1,\dots,N}$ for an observable $g\in L^{2}(X)$ and a set of learned basis functions $\{\phi_{j}\}_{j=1}^{N}$. In more general terms, we define $\mathcal{P}_{\{\phi_{j}\}_{j=1}^{N}}:L^{2}(X)\to\mathbb{R}^{N}$ via
@@
We shall sometimes omit the subscript of $\mathcal{P}$.
@@
The action of the transfer/Koopman operator is represented by a linear map implemented in block $\mathcal{G}$. The coefficients $\langle g,\phi_{j}\rangle$, $j=1,\ldots,N$ in the latent space are fed into $\mathcal{G}$, which learns the action of $\hat{\mathcal{L}}$ in the latent space. That is, the block $\mathcal{G}$ consists of a learnable linear map capable of transforming the coefficients of $g$ when represented in the basis $\{\phi_{j}\}_{j=1}^{N}$ into the corresponding coefficients of ^$\hat{\mathcal{L}}(g)$ in the same basis.
Reconstruction ($\mathcal{R}$)
The architecture employs a reconstruction block that converts the latent coefficients produced by the linear map $\mathcal{G}$ into the approximation of the transformed observable
@@
Because the same functional expression can reconstruct any element of the span of the learned basis, we keep the generic definition of $\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}\colon\mathbb{R}^{N}\to L^{2}(X)$ as
(3) $\mathcal{R}_{\{\phi_{j}\}_{j=1}^{N}}(a_{1},\dots,a_{N})=\sum_{j=1}^{N}a_{j}\,\phi_{j}.$
To simplify the notation, we will also sometimes omit the subscript in the sequel. The coefficients serving as input to $\mathcal{R}$ will always be clear from the context (i.e. which maps are composed with $\mathcal{R}$).
One of the key features of SABON is that it simultaneously optimises the learned subspace and the operator action. This is performed through a suitable loss function comprised of two components (in practice these are calculated with their discrete counterparts using evaluations at the sample points $\{x_{i}\}_{i=1}^{n}$):
@@
    The composition $\mathcal{R}\circ\mathcal{G}\circ\mathcal{P}$, which approximates the projected action of $\mathcal{L}$ on $V$, has the loss term:
    (4) $\mathcal{E}_{1}(\mathcal{L},\mathcal{G},\mathcal{P},\mathcal{R},g)\;\coloneq\;\frac{\|\mathcal{L}(g)-\mathcal{R}\circ\mathcal{G}\circ\mathcal{P}(g)\|_{2}}{\|\mathcal{L}(g)\|_{2}}.$
@@
    A sparsity penalty encourages the learned basis functions to have local support:
    (5) $\mathcal{E}_{2}(\{\phi_{j}\}_{j=1}^{N})=\frac{1}{N}\sum_{j=1}^{N}\|\phi_{j}\|_{L^{1}}.$
@@
(6) $J(\mathcal{L},\mathcal{G},\mathcal{P},\mathcal{R},g,\{\phi_{j}\}_{j=1}^{N})=\beta_{1}\mathcal{E}_{1}(\mathcal{L},\mathcal{G},\mathcal{P},\mathcal{R},g)+\beta_{2}\mathcal{E}_{2}(\{\phi_{j}\}_{j=1}^{N}),$
where $\beta_{1},\beta_{2}\geq 0$.
We also explored supplementary projection penalties aimed at reconstructing observables after their projection onto the subspace. The first variant penalised deviations in reconstructing the original observable:
$\mathcal{E}_{p_{1}}(\mathcal{P},\mathcal{R},g)\;\coloneq\;\frac{\|\mathrm{Id}(g)-\mathcal{R}\circ\mathcal{P}(g)\|_{2}}{\|\mathrm{Id}(g)\|_{2}},$
while the second variant addressed the reconstruction of $\mathcal{L}(g)$ after projection onto the learned subspace:
$\mathcal{E}_{p_{2}}(\mathcal{P},\mathcal{R},\mathcal{L},g):=\frac{\|\mathrm{Id}(\mathcal{L}(g))-\mathcal{R}\circ\mathcal{P}(\mathcal{L}(g))\|_{2}}{\|\mathrm{Id}(g)\|_{2}}$
where $\mathrm{Id}$ is the identity operator on $L^{2}(X)$. In practice, we found that the first penalty sometimes adversely affected the invariance property central to our goals — eigenpair reconstruction and subspace invariance. Conversely, the second penalty demonstrated a neutral impact, presumably due to its redundancy with $\mathcal{E}_{1}$, since effective operator approximation implicitly requires accurate reconstruction after projection.
Finally, we discuss the activation functions employed in the proposed architecture. For the $\mathcal{G}$ network, which is learning the action of a linear operator (transfer or Koopman), we selected a linear activation function, namely the identity function, and set the bias term to zero. The activation functions for the encoder block ($\mathcal{E}$) can be chosen among the usual candidates, and we had the best results with ReLU in our experiments.
The combined architecture and loss in SABON are structured so that the learned basis $\{\phi_{j}\}_{j=1}^{N}$ (i) is approximately orthogonal, (ii) spans a subspace $V$ that is very close to invariant under the Perron–Frobenius/Koopman operator, and (iii) has a tendency toward local support, guided by the sparsity term $\mathcal{E}_{2}$ in the loss function.
Further, $\mathcal{E}_{1}$ simultaneously optimises for a good finite–dimensional approximation of the operator action and enforces that $\{\phi_{j}\}_{j=1}^{N}$ spans a space $V$ well suited for representing $\mathcal{L}(g)$. A final rescaling step yields an orthonormal basis.